# Efficient Event-based Semantic Segmentation with Spike-driven Lightweight Transformer-based Networks

### v1

@@ Abstract
Event-based semantic segmentation has great potential in autonomous driving and robotics due to the advantages of event cameras, such as high dynamic range, low latency, and low power cost. Unfortunately, current artificial neural network (ANN)-based segmentation methods suffer from high computational demands, the requirements for image frames, and massive energy consumption, limiting their efficiency and application on resource-constrained edge/mobile platforms. To address these problems, we introduce SLTNet, a spike-driven lightweight transformer-based network designed for event-based semantic segmentation. Specifically, SLTNet is built on efficient spike-driven convolution blocks (SCBs) to extract rich semantic features while reducing the model’s parameters. Then, to enhance the long-range contextural feature interaction, we propose novel spike-driven transformer blocks (STBs) with binary mask operations. Based on these basic blocks, SLTNet employs a high-efficiency single-branch architecture while maintaining the low energy consumption of the Spiking Neural Network (SNN). Finally, extensive experiments on DDD17 and DSEC-Semantic datasets demonstrate that SLTNet outperforms state-of-the-art (SOTA) SNN-based methods by at least 7.30% and 3.30% mIoU, respectively, with extremely 5.48$\times$ lower energy consumption and 1.14$\times$ faster inference speed.
@@ I INTRODUCTION
Semantic segmentation [1] aims to divide visual data into distinct regions with clear semantic properties, thereby enabling a deep understanding and analysis of the scene. This technology plays a crucial role in various cutting-edge fields such as autonomous driving, security monitoring [2], and robotics [3], forming the foundational component for environmental perception capabilities of intelligent robotics [4]. However, with the increasing accuracy and efficiency demands for robotics in high-dynamic, complex scenarios, traditional image-based semantic segmentation methods show significant performance degradation when processing high-dynamic or variable lighting scenes [3].
@@
Event cameras, leveraging their innovative biomimetic event generation characteristics, have inspired considerable research interests in computer vision and robotics communities [5]. Event cameras [6] excel at detecting variations in light intensity at the pixel level with remarkable microsecond resolution. They asynchronously produce an event stream that consists of time stamp $t$, spatial coordinates $(x,y)$, and polarity (+,)(+,-)( + , - ). The event-driven nature of these cameras shows several superior advantages, including HDR perception, low latency, low energy cost, and enhanced sparsity, showing promising solutions to computationally intensive semantic segmentation. Although research on event-based semantic segmentation is still in its infancy, several works have been proposed. Ev-SegNet [7] develops the first baseline using solely events but requires a substantial computational load. Following this, DTL [8], Evdistill [9], and ESS [10] attempt to compensate for the lack of visual details in event data through transfer learning or knowledge distillation technologies from image processing branches. However, existing methods either are still computationally expensive or require auxiliary images. Therefore, a critical challenge is how to design an efficient and robust event-based segmentation model with lower computational and energy costs, sufficient event feature extraction ability, and without requirement for auxiliary modality.
@@
With the development of biological computing, Spiking Neural Network (SNNs) is a potential solution inspired by mechanisms of information generation and transmission in the brain, which can convey dynamically generated binary spikes across spatial and temporal dimensions [11]. Thus, SNNs demonstrate excellent advantages in computation efficiency, temporal memory capacity, and biological interpretability. The basic building block of SNNs is the spiking neuron, which shares a highly similar generation mechanism with the event cameras, making SNNs theoretically well-suited to work with such cameras. Encouraged by these advantages, Kim et al. [12] attempt to integrate SNNs with classical ANN-based segmentation networks, but the performance is poor due to the plain feature extraction module. This indicates that combining the SNNs with event segmentation still requires further exploration.
@@
To address the aforementioned issues, we propose SLTNet, a Spike-driven Lightweight Transformer-based semantic segmentation network that utilizes events only to achieve energy-efficient and robust performance. SLTNet is a hierarchical single-branch SNN-based method equipped with an encoder-decoder structure. It can extract both detail and semantic information while maintaining high efficiency and low energy consumption. Specifically, as shown in Fig. 3, SLTNet contains four stages in the encoder part, which includes three novel Spike-driven Convlution Blocks (SCBs) that capture fine-grained textural information, and two newly designed Spike-driven Transformer Blocks (STBs) that are responsible for extracting long-range contextual features in the last stage. Then, to further reduce the computational cost, we propose a Spiking Lightweight Dilated module (Spike-LD) that serves as the basic blocks of our conv-based SCBs and decoders. Finally, extensive experimental results conducted on two public event segmentation datasets demonstrate that the SLTNet outperforms competitive SOTA methods by at least 7.30% and 3.30% mIoU.
@@
  • To enhance feature interaction capabilities, we propose novel Spike-driven Transformer Blocks, named STBs, by leveraging the re-parameterization convolution technique and spike-driven multi-head self-attention mechanism to achieve long-range contextural feature interaction with only floating-point accumulation operations.
  • Extensive experimental results show that SLTNet outperforms SNN-based methods by at least 7.30% and 3.30% mIoU with much lower computation cost.
II EFFICIENT BUILDING BLOCKS
II-A Evolutionary Spiking Neuron Module
@@
where the $U^{t}_{l}$ is the membrane potential of neurons in the $l^{th}$ layer at time step $t$, $W^{t}_{l}$ is synaptic weight matrix between layers $l-1$ and $l$, $S^{t}_{l}$ is the output of current layer, and $\tau$ is membrane time constant. The LIF module assumes that neurons accumulate membrane potential like a capacitor. They receive the leaky membrane potential of the previous time step $\tau U^{t-1}_{l}$ and calculate current potential $W^{t}_{l}X^{t}_{l-1}$. When the accumulation reaches a certain threshold $U_{th}$, the neuron fires, generating a spike $S^{t}_{l}$ using Heaviside step function $H(\cdot)$; otherwise, it remains zero. After firing, the membrane potential rebound to $U_{reset}$.
@@
II-B Spiking Lightweight Dilated Module
@@
II-C Spike-driven Transformer Block
@@
III OVERALL ARCHITECTURE OF SLTNET
To reduce computational load while maintaining semantic information from events, we design a Spike-driven Lightweight Transformer-based Network, called SLTNet, for event-based semantic segmentation based on the tailored efficient Spike-LD modules and STBs.
III-A Event Representation
@@
III-B Spike-driven Encoder/Decoder of SLTNet
@@
III-C Segmentation Loss Design
@@
IV EXPERIMENTS AND RESULTS
IV-A Experimental Datasets
@@ IV-C Comparision with State-of-the-art Methods
DDD17 Ev-SegNet Event ANN 23.75 44.17 20.67 54.81
Evdistill Event+Frame ANN 58.64 65.08 2.96 58.02
ESS Event+Frame ANN 6.69 - - 61.37
Spiking-DeepLab Event SNN 15.89 4.05 100 44.63
Spiking-FCN Event SNN 18.64 35.23 72.73 42.87
@@
DSEC-Semantic Ev-SegNet Event ANN 23.75 - - 51.76
ESS Event+Frame ANN 6.69 - - 51.57
Spiking-DeepLab Event SNN 15.89 14.02 100 44.58
Spiking-FCN Event SNN 18.64 126.46 17.78 38.52
@@
IV-C1 Performance on DDD17
Detailed results are given in Tab. I. In the field of SNNs, our SLTNet achieves the highest segmentation performance while maintaining the smallest number of parameters, minimal floating-point computations, and the fastest inference speed. Specifically, we achieve 7.30% and 9.06% higher mIoU than Spiking-DeepLab and Spiking-FCN, respectively. Meanwhile, it has 38.76$\times$ fewer parameters, 2.07$\times$ less GFLOPs, and 1.14$\times$ faster FPS than comparative SNN methods.
Next, we compare SLTNet to other ANN methods. Though our SLTNet slightly lags behind Ev-SegNet, it requires 57.93$\times$ fewer parameters, 22.54$\times$ less FLOPs, and achieves 5.53$\times$ faster FPS. Although there is a performance gap in mIoU compared to Evdistill and ESS, SLTNet achieves these results without using any image information and with significantly lower energy consumption and parameters. The energy consumption will be discussed in detail later. We present qualitative results in Fig. 4 to provide a more intuitive comparison of different models. Our proposed method achieves fine-grained segmentation compared to competition models, which demonstrates the effectiveness of our design.
IV-C2 Performance on DSEC-Semantic
As shown in Table. I, our model achieves the best segmentation result within the SNN field, with a performance of 47.91% mIoU. Specifically, it outperforms Spiking-DeepLab by 3.33% with 9.51$\times$ fewer parameters and Spiking-FCN by 9.39% with 11.16$\times$ fewer parameters. Though our model still falls behind ANN-based methods, the gap between them has significantly narrowed compared to the DDD17 dataset as the spatial resolution of the event data increases. At the same time, SLTNet achieves fewer parameters (1.67 M) and FLOPs (6.97 G) across all ANN and SNN methods, which demonstrates superior performance in computation efficiency and process speed, achieves promising 114.29 FPS.
@@ IV-C2 Performance on DSEC-Semantic
IV-D Power and Computation Cost Analysis
@@
Table II compares energy consumption with other models. To maintain consistency with previous work [12], we set $E_{MAC}=4.6pJ$ and $E_{AC}=0.9pJ$ during calculation. Compared to ANN-based models, our proposed SLTNet excels in energy efficiency. Specifically, Ev-SegNet, Evdistill, and ESS consume 30.20$\times$, 96.31$\times$, and 37.90$\times$ more energy than our model, respectively. Although Spiking-DeepLab and Spiking-FCN also utilize spiking architectures, their straightforward components and design result in 5.48$\times$ and 12.00$\times$ higher power consumption.
IV-E Ablation Studies
IV-E1 Spike-driven Multi-head Self-attention
@@ IV-E2 Shortcut and Architecture
The ablation parts on the network architecture have almost no impact on the number of parameters, effectively eliminating the effect of parameter space on performance. Shortcut. There are three shortcuts in SNNs, where we use the membrane shortcut (MS) that connects float membrane potential in different layers. The other two are the Spike-Element-Wise (SEW) [19] Shortcut, which fuses spikes across layers, and the Vanilla Shortcut (VS) [18], which directly connects spikes to membrane potential. As shown in Table IV, in our architecture, our MS performs the best, with VS reducing the segmentation accuracy by 2.90% and SEW shortcut by 19.38%. The poor performance of SEW shortcut indicates that using spikes alone without retaining information about the membrane potential can significantly impair performance on tasks requiring detailed information. Architecture. When removing the FE module in the skip connection, the mIoU is reduced by 1.96%, demonstrating its effectiveness in enhancing feature expression. Furthermore, we change the network to fully SCBs without STBs. Performance dropped by 3.46% significantly, which proves that transformer and convolution blocks can achieve complementarity. Then, we remove the skip connection from stages 1-3 as described in Sec. III-B. It leads to a performance loss of 2.27%, showing that the feature map from the early stage can effectively complement segmentation information.

### v2

# SLTNet: Efficient Event-based Semantic Segmentation with Spike-driven Lightweight Transformer-based Networks

@@ Abstract
Event-based semantic segmentation has great potential in autonomous driving and robotics due to the advantages of event cameras, such as high dynamic range, low latency, and low power cost. Unfortunately, current artificial neural network (ANN)-based segmentation methods suffer from high computational demands, the requirements for image frames, and massive energy consumption, limiting their efficiency and application on resource-constrained edge/mobile platforms. To address these problems, we introduce SLTNet, a spike-driven lightweight transformer-based network designed for event-based semantic segmentation. Specifically, SLTNet is built on efficient spike-driven convolution blocks (SCBs) to extract rich semantic features while reducing the model’s parameters. Then, to enhance the long-range contextural feature interaction, we propose novel spike-driven transformer blocks (STBs) with binary mask operations. Based on these basic blocks, SLTNet employs a high-efficiency single-branch architecture while maintaining the low energy consumption of the Spiking Neural Network (SNN). Finally, extensive experiments on DDD17 and DSEC-Semantic datasets demonstrate that SLTNet outperforms state-of-the-art (SOTA) SNN-based methods by at most 9.06% and 9.39% mIoU, respectively, with extremely 4.58$\times$ lower energy consumption and 114 FPS inference speed. Our code is open-sourced and available at https://github.com/longxianlei/SLTNet-v1.0.
@@ I Introduction
Semantic segmentation [1] aims to divide visual data into distinct regions with clear semantic properties, thereby enabling a deep understanding and analysis of the scene. This technology plays a crucial role in various cutting-edge fields such as autonomous driving  [2], security monitoring [3], and robotics [4], forming the foundational component for environmental perception capabilities of intelligent robotics [5]. However, with the increasing efficiency demands for robotics in complex scenarios, traditional image-based semantic segmentation methods show significant performance degradation when processing high-dynamic or variable lighting scenes. This challenge has driven the development of event camera-based applications.
@@
Event cameras, leveraging their innovative biomimetic event generation characteristics, have inspired considerable research interests in computer vision and robotics communities [6]. Event cameras [7] excel at detecting variations in light intensity at the pixel level with remarkable microsecond resolution. They asynchronously produce an event stream that consists of time stamp $t$, spatial coordinates $(x,y)$, and polarity (+,)(+,-)( + , - ). The event-driven nature of these cameras shows several superior advantages, including high dynamic range (HDR) perception, low latency, and low energy cost, showing promising solutions to computationally intensive semantic segmentation [8, 9]. Although research on event-based semantic segmentation is still in its infancy, several works have been proposed. Ev-SegNet [10] develops the first baseline using solely events but requires a substantial computational load. Following this, DTL [11], Evdistill [12], and ESS [13] attempt to compensate for the lack of visual details in event data through transfer learning or knowledge distillation technologies from image processing branches. However, existing methods remain computationally expensive or rely on auxiliary images [14]. Therefore, a critical challenge is to design an efficient and robust event-based segmentation model that minimizes computational and energy costs while boosting effective event feature extraction ability and reducing model complexity.
@@
With the development of biological computing, Spiking Neural Network (SNNs) is a potential solution inspired by mechanisms of information generation and transmission in the brain, which can convey dynamically generated binary spikes across spatial and temporal dimensions [15]. Thus, SNNs demonstrate excellent advantages in computation efficiency, temporal memory capacity, and biological interpretability. The basic building block of SNNs is the spiking neuron, which shares a highly similar generation mechanism with the event cameras, making SNNs theoretically well-suited to work with such cameras. Encouraged by these advantages, some works [16, 17, 18, 19] attempt to integrate SNNs with classical ANN-based segmentation networks, but the performance is poor due to the plain feature extraction module. This indicates that combining the SNNs with event segmentation still requires further exploration.
@@
To address these issues, we propose SLTNet, a Spike-driven Lightweight Transformer-based semantic segmentation network that utilizes events only to achieve energy-efficient and robust performance. SLTNet is a hierarchical single-branch SNN-based method equipped with an encoder-decoder structure. It can extract both detailed and semantic information from only event data while maintaining high efficiency and low energy consumption. Specifically, SLTNet contains four stages in the encoder part, which includes three novel Spike-driven Convlution Blocks (SCBs) that capture fine-grained textural information, and two newly designed Spike-driven Transformer Blocks (STBs) that are responsible for extracting long-range contextual features in the last stage. Then, to further reduce the computational cost, we propose a Spiking Lightweight Dilated module (Spike-LD) that serves as the basic blocks of our conv-based SCBs and decoders. Finally, extensive experimental results conducted on two recognized public event segmentation datasets demonstrate that the SLTNet outperforms SNN-based SOTA methods by a large margin while maintaining extremely lower energy consumptions with faster inference speed.
@@
  • To enhance feature interaction capabilities, we propose novel Spike-driven Transformer Blocks (STBs), by leveraging the re-parameterization convolution technique and spike-driven multi-head self-attention mechanism to achieve long-range contextual feature interaction with only floating-point accumulation operations.
@@
  • Extensive experimental results show that SLTNet outperforms SNN-based methods by at most 9.06% and 9.39% mIoU with much lower computation cost and 114 FPS inference speed.
II Related Works
II-A Event-based Semantic Segmentation
Event cameras [7] feature a unique event-driven mechanism that makes them highly sensitive to moving objects. This high sensitivity presents a challenge in developing novel algorithms to extract meaningful information from event data effectively. With the development in deep learning and the availability of large-scale labeled event datasets, event camera-based applications have become a prominent research area, including their application in semantic segmentation [9].
Ev-SegNet [10] introduces the first baseline for semantic segmentation using only event data, leveraging an Xception-like model and an automatic labeling method to generate approximate semantic annotations. Although semantic segmentation is well-suited for image data, this work demonstrates its feasibility with event data. However, the sparse nature of event streams and the limited availability of high-quality labeled datasets hinder the development of event-based methods to surpass image-based approaches. To address these, recent efforts focus on transferring knowledge from images to enhance performance. E.g., EvDistill [12] employs knowledge distillation to incorporate features from a teacher network trained on large-scale labeled image data, achieving remarkable performance on unlabeled and unpaired event data. Similarly, ESS [13] utilizes unsupervised domain adaptation to transfer semantic segmentation tasks from labeled image datasets to unlabeled event data.
However, existing studies adopt processing paradigms originally designed for images, neglecting the inherent spatiotemporal relationships within event streams. As a result, event-based semantic segmentation algorithms remain underdeveloped, highlighting the need for our proposed approach.
II-B SNN-based Semantic Segmentation
To tackle the above problems, researchers resort to the SNN for its superior performance in energy efficiency, temporal memory capacity, and biological interpretability [20]. Kim et al. [16] and EvSegSNN [17] incorporate spiking neurons into classical FCN, DeepLab, and U-Net segmentation architectures, but these models showed unsatisfactory performance. Then, MFS-EFS [19] proposes a spiking context-guided architecture with biologically inspired feature extraction blocks, enabling unified semantic segmentation across frame and event-based vision modalities through direct spike-driven training. To enhance performance, SCGNet [18] introduces a multi-scale fully-spiking segmentation architecture with depthwise-separable residual blocks, pioneering surrogate gradient-based training of deep SNNs for semantic segmentation. It also theoretically mitigates spiking noise and degradation through dynamical isometry principles.
However, existing approaches still exhibit critical deficiencies in both energy efficiency and computational performance. Designing an efficient and compact architecture remains a challenge. This work aims to achieve an optimal balance by introducing a novel SNN architecture specifically designed for event stream segmentation.
III Efficient Building Blocks of SLTNet
III-A Evolutionary Spiking Neuron Module
@@
where the $U^{t}_{l}$ is the membrane potential of neurons in the $l^{th}$ layer at time step $t$, $W^{t}_{l}$ is synaptic weight matrix between layers $l-1$ and $l$, $S^{t}_{l}$ is the output of current layer, and $\tau$ is membrane time constant. The LIF module receives the leaky membrane potential of the previous time step $\tau U^{t-1}_{l}$ and calculates current potential $W^{t}_{l}X^{t}_{l-1}$. When the accumulation reaches threshold $U_{th}$, the neuron fires, generating a spike $S^{t}_{l}$ using Heaviside step function $H(\cdot)$; otherwise, it remains zero. After firing, the membrane potential rebounds to $U_{reset}$.
@@
III-B Spiking Lightweight Dilated Module
@@
III-C Spike-driven Transformer Block
@@
IV Overall Architecture of SLTNet
To reduce computational load while maintaining semantic information from events, based on the tailored efficient Spike-LD modules and STBs building blocks introduced in Sec. III, we design SLTNet for event semantic segmentation.
IV-A Event Representation
@@
IV-B Spike-driven Encoder/Decoder of SLTNet
@@
IV-C Segmentation Loss Design
@@
V Experiments and Results
V-A Experimental Datasets
@@ V-B Implementation Details
DDD17 Ev-SegNet [10] Event ANN 23.75 44.17 20.67 54.81
Evdistill [12] Event+Frame ANN 58.64 65.08 2.96 58.02
ESS [13] Event+Frame ANN 6.69 - - 61.37
Spiking-DeepLab [16] Event SNN 15.89 4.05 100 44.63
Spiking-FCN [16] Event SNN 18.64 35.23 72.73 42.87
EvSegSNN [17] Event SNN 8.55 - - 45.54
SCGNet [18] Event SNN 0.49 - - 49.30
@@
DSEC-Semantic Ev-SegNet [10] Event ANN 23.75 - - 51.76
ESS [13] Event+Frame ANN 6.69 - - 51.57
Spiking-DeepLab [16] Event SNN 15.89 14.02 100 44.58
Spiking-FCN [16] Event SNN 18.64 126.46 17.78 38.52
MFS-EFS [19] Event SNN 1.11 28.88 - 46.10
@@
V-C Comparision with State-of-the-art Methods
V-C1 Performance on DDD17
Detailed results are given in Tab. I. In the field of SNNs, our SLTNet achieves the highest segmentation performance while maintaining the smallest number of parameters, minimal floating-point computations, and the fastest inference speed. Specifically, we achieve 7.30%, 9.06%, 6.39% and 2.63% higher mIoU than Spiking-DeepLab [16], Spiking-FCN [16], EvSegSNN [10], and SCGNet [18] respectively. Meanwhile, it has fewer parameters, 2.07$\times$ less FLOPs, and 1.14$\times$ faster FPS than comparative SNN methods.
Next, we compare SLTNet to other ANN methods. Though our SLTNet slightly lags behind Ev-SegNet, it requires 57.93$\times$ fewer parameters, 22.54$\times$ less FLOPs, and achieves 5.53$\times$ faster FPS. Although there is a performance gap in mIoU compared to Evdistill and ESS, SLTNet achieves these results without using any image information and with significantly lower energy consumption and parameters. The energy consumption will be discussed in detail later.
V-C2 Performance on DSEC-Semantic
As shown in Table. I, our model achieves the best segmentation result within the SNN field, with a performance of 47.91% mIoU. Specifically, it outperforms Spiking-DeepLab by 3.33% with 9.51$\times$ fewer parameters and Spiking-FCN by 9.39% with 11.16$\times$ fewer parameters. Furthermore, our method shows a 1.81% performance enhancement over the latest MFS-EFS while operating with 4.14$\times$ fewer FLOPs. Though our model still falls behind ANN-based methods, the gap between them has significantly narrowed compared to the DDD17 dataset as the spatial resolution of the event data increases. At the same time, SLTNet achieves fewer parameters (1.67 M) and FLOPs (6.97 G) across all ANN and SNN methods, which demonstrates superior performance in computation efficiency and process speed, achieves promising 114.29 FPS.
We also present qualitative results in Fig. 4 and Fig. 5 to provide a more intuitive comparison of different models. It clearly shows that our proposed SLTNet achieves fine-grained segmentation compared to competition models, demonstrating the robustness and effectiveness of our model across different datasets.
@@ V-C2 Performance on DSEC-Semantic
Figure 5: Segmentation results on the DSEC-Semantic dataset. The left are event frames and corresponding image frames. The rightmost are ground truths.
@@
*: The energy consumption reported in SCGNet’s [18] original paper is 6.50 $mJ$. For consistency, we present this value without modification.
V-D Power and Computation Cost Analysis
@@
Table II compares energy consumption with other models. To maintain consistency with previous work [16], we set $E_{MAC}=4.6pJ$ and $E_{ACC}=0.9pJ$ during calculation. Compared to ANN-based models, our proposed SLTNet excels in energy efficiency. Specifically, Ev-SegNet, Evdistill, and ESS consume 30.20$\times$, 96.31$\times$, and 37.90$\times$ more energy than our model, respectively. Although Spiking-DeepLab, Spiking-FCN and SCGNet also utilize spiking architectures, their straightforward components and design result in 5.48$\times$, 12.00$\times$ and 4.58$\times$ higher power consumption.
V-E Ablation Studies
V-E1 Spike-driven Multi-head Self-attention
@@ V-E2 Shortcut and Architecture
The ablation parts on the network architecture have almost no impact on the number of parameters, effectively eliminating the effect of parameter space on performance.
Shortcut. There are three shortcuts in SNNs, where we use the membrane shortcut (MS) that connects float membrane potential in different layers. The other two are the Spike-Element-Wise (SEW) [27] Shortcut, which fuses spikes across layers, and the Vanilla Shortcut (VS) [26], which directly connects spikes to membrane potential. As shown in Table IV, in our architecture, our MS performs the best, with VS reducing the segmentation accuracy by 2.90% and SEW shortcut by 19.38%. The poor performance of SEW shortcut indicates that using spikes alone without retaining information about the membrane potential can significantly impair performance on tasks requiring detailed information.
Architecture. When removing the FE module in the skip connection, the mIoU is reduced by 1.96%, demonstrating its effectiveness in enhancing feature expression. Furthermore, we change the network to fully SCBs without STBs. Performance dropped by 3.46% significantly, which proves that transformer and convolution blocks can achieve complementarity. Then, we remove the skip connection from stages 1-3 as described in Sec. IV-B. It leads to a performance loss of 2.27%, showing that the feature map from the early stage can effectively complement segmentation information.